OpenStudy (anonymous):

The point P(8, 6) is on the terminal side of θ. Evaluate sin θ.

OpenStudy (anonymous):

 sin θ is = opposite/adj. Can you find it?

OpenStudy (anonymous):

no, I cannot

OpenStudy (anonymous):

what is the opposite side of θ?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

yes and what is beside θ

OpenStudy (anonymous):

8

OpenStudy (anonymous):

so if sin θ = opposite/ adjacent(beside) what is the ratio?

OpenStudy (anonymous):

sin O = 6/8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

those arent my options though, not even 3/4

OpenStudy (anonymous):

5/4, 4/3, 4/5, 3/5

OpenStudy (anonymous):

are my options

OpenStudy (anonymous):

hmmm.... are you sure that yea I forgot lol... my bad sin θ = opp/ hypotenuse lol

OpenStudy (anonymous):

you use Pythagorean thearom to solve hypo

OpenStudy (anonymous):

8^2 +6^2 = c^2

OpenStudy (anonymous):

solve c

OpenStudy (anonymous):

OHHHHHH, GOTCHA, EASY PEASY

OpenStudy (anonymous):

so its 10

OpenStudy (anonymous):

you should get like 3/5...so yea that's it.